# K3 surfaces with real or complex multiplication

### v2

@@ 10.1. Hodge structures of K3 type
The polarization is negative definite on $U_{0}:=(U\otimes_{{\mathbf{Z}}}\mathbf{R})\cap U^{1,1}_{X}$ and thus has signature $(2,r-2)$ on $U_{\mathbf{R}}=U_{\mathbf{Z}}\otimes_{\mathbf{Z}}\mathbf{R}$.
A simple K3 type Hodge structure is a member of an $r-2$-dimensional family of K3 type Hodge structures, the general one again being simple, which is parametrized by the choice of an $\omega\in U\otimes_{\mathbf{Z}}\mathbf{C}$ with $q(\omega,\omega)=0$ and $q(\omega,\bar{\omega})>0$; the Hodge structure $U_{\omega}$ determined by $\omega$ has $U^{2,0}_{\omega}=\mathbf{C}\omega$.
Any rational polarized Hodge structure $(U,q)$ of K3 type can be realized geometrically as a Hodge substructure of $H^{2}(KS(U),{\mathbf{Q}})$ where $KS(U)$ is the Kuga Satake abelian variety $KS(U)$ of $U$, which is defined in terms of the Clifford algebra of $(U,q)$. Combined with Lefschetz theorems, it follows that these Hodge structures also appear in the second cohomology group of certain projective surfaces. In this paper, however, we will be mainly interested in such Hodge structures that arise from transcendental lattices in the second cohomology groups of hyperkähler manifolds.
@@ 10.2. Endomorphism algebras of K3 type Hodge structures
$A_{U}\,=\,{\rm End}_{\rm Hdg}(U)\,=\,\{f\in{\rm End}(U):\;f(U^{p,q})\subset U^{p,q}\}\leavevmode\nobreak\ .$
By considering the action of an $f\in A_{U}$ on $U^{2,0}\simeq\mathbf{C}$ and the polarization on $U$, Zarhin [Z 83] showed that $A_{U}$ is either a CM field or a totally real field. The adjoint of $a\in A_{U}$ for the polarization is the complex conjugate of $a$:
$q(ax,y)\,=\,q(x,\bar{a}y)\leavevmode\nobreak\ .$
This adjoint property is equivalent to the existence of an isometry of quadratic forms $U\simeq\mathrm{T}(W)$ for some quadratic (or hermitian) form $W$ over the RM (or CM) field $A_{U}$, see Lemma 4.1. In the RM case, the form $W$ must be such that $W_{\sigma}$ has signature $(2,m-2)$ where $m=\dim W$ and $m\geq 3$ for one embedding and thus $W_{\tau}$ is negative definite for all other embeddings. For totally real fields, there is the following restriction, but for CM fields there is none.
@@
The results in Section 9 characterize the fields $E$, totally real or CM, such that $(U,q)=\mathrm{T}(W)$. The results in Sections 12, 14.1 and 14.2 apply these criteria to specific quadratic forms that arise from Hodge substructures $T\subset H^{2}(X,{\mathbf{Z}})$ of K3 type, where $X$ is a hyperkähler manifold.
@@ 10.3. $K3$ surfaces
We recall the basic facts on K3 surfaces needed for the application of the results obtained in the previous sections.
@@
$H^{2}(X,{\mathbf{Z}})\otimes_{{\mathbf{Z}}}\mathbf{C}\,=\,H^{2,0}(X)\oplus H^{1,1}(X)\oplus H^{0,2}(X),\qquad H^{0,2}(X)\,=\,\overline{H^{2,0}(X)}\leavevmode\nobreak\ ,$
@@ 10.4. Hyperkähler manifolds
A hyperkähler (HK) manifold, also called irreducible holomorphic symplectic manifold (IHSM), is a simply connected compact complex Kähler manifold $X$ with trivial canonical bundle. A 2-dimensional HK manifold is a K3 surface. Currently we know of four higher dimensional families of such manifolds, listed in Table 1:
@@
The second integral cohomology group $H^{2}(X,{\mathbf{Z}})$ of a HK manifold shares many of the properties of the one of a K3 surface. It is a free ${\mathbf{Z}}$-module of rank $r=b_{2}(X)$ with an even quadratic form $q$ (not unimodular in general), the Beauville-Bogomolov-Fujiki-form (or BBF-form), of signature $(3,r-3)$. The Hodge decomposition has
$H^{2,0}(X)\,=\,\mathbf{C}\omega\;\;\;\text{ and }\;\;\;\mathrm{Pic}(X)=\omega^{\perp}\cap H^{2}(X,{\mathbf{Z}})\leavevmode\nobreak\ ,$
where $\omega\in H^{2}(X,\mathbf{C})$ is (the class of) a nowhere zero holomorphic two form. A HK manifold $X$ is projective if and only if $\mathrm{Pic}(X)$ contains an element $D$ with $q(D,D)>0$. The BBF-form defines a polarization $q$ on the Hodge structure $T_{X}:=\mathrm{Pic}(X)^{\perp}$ and $(T_{X},q)$ is a simple Hodge structure of K3 type. We write $L_{r,n}$ for the lattice $(H^{2}(X,{\mathbf{Z}}),q)$ (which only depends $r=b_{2}(X)$ and $n=\dim X$):
$L_{r,n}\,\simeq\,(H^{2}(X,{\mathbf{Z}}),q),\quad V_{r,n}\,:=\,L_{r,n}\otimes_{\mathbf{Z}}{\mathbf{Q}}\,\simeq\,(H^{2}(X,{\mathbf{Q}}),q)\leavevmode\nobreak\ .$
@@ 10.5. Surjectivity of the period map for HK manifolds
The main result from complex geometry that we need to establish the existence of HK manifolds with RM or CM is the surjectivity of the period map. This is the fact ([H 99]), that if $L_{r,n}=(V_{\mathbf{Z}},q_{V})$ is a lattice which is isometric to $(H^{2}(X,{\mathbf{Z}}),q)$ for a HK manifold $X$ and given an $\omega\in V_{\mathbf{Z}}\otimes_{{\mathbf{Z}}}\mathbf{C}$ with $q_{V}(\omega,\omega)=0$ and $q_{V}(\omega,\overline{\omega})>0$, then there exists a HK manifold $X_{\omega}$ with an isometry $(H^{2}(X_{\omega},{\mathbf{Z}}),q)\simeq(V_{\mathbf{Z}},q)$ that maps $H^{2,0}(X_{\omega})$ to $V^{2,0}:=\mathbf{C}\omega$. Moreover, if there there exists an $h\in V_{\mathbf{Z}}\cap\omega^{\perp}$ with $q_{V}(h,h)>0$, then $X_{\omega}$ is a projective HK manifold. From the Lefschetz $(1,1)$-theorem it follows that
[…]
@@
For K3 surfaces there is the Torelli theorem which asserts that $X_{\omega}$ is uniquely determined by $\omega$ up to isomorphism. This does not hold in general for HK manifolds, but it is still true that fibers of the period map are zero dimensional.
@@ 11.1. The condition (C)
Let $E$ be a totally real or CM number field of degree $d$. Let $U=(U,q)$ be a quadratic form over ${\mathbf{Q}}$ of signature $(2,r-2)$.
@@ 11.2. Geometric realizations
Let $(U,q)$ a quadratic form over ${\mathbf{Q}}$ which is a direct summand of the quadratic form $V_{r,n}\simeq(H^{2}(X,{\mathbf{Q}}),q)$ as in Table 1; notice that this quadratic form only depends $r=b_{2}(X)$ and $n=\dim X$. Let $E$ be a field such that $U$ admits an $E$-module structure satisfying condition (C). Then the choice of a suitable, general, $\omega$ in a suitable eigenspace of the $E$-action on $U_{\mathbf{C}}$ determines a Hodge structure on $V_{r,n}$ and thus on $L_{r,n}\subset V_{r,n}$. By the surjectivity of the period map this determines a finite number of HK manifolds $X$ (and $X$ is unique if $r=22$) with a Hodge isometry $U\simeq T_{X,{\mathbf{Q}}}$ and $E\subset A_{X}$; equality can often be established with dimension arguments.
The results obtained in the previous sections on quadratic forms allow us to find examples of such $E$ and $U$. With this in mind, the following result provides the key geometric ingredient to prove Theorems A and B.
@@
Let $V_{r,n}=L_{r,n}\otimes_{\mathbf{Z}}{\mathbf{Q}}$ where $L_{r,n}$ is a lattice from Table 1. Let $E$ be a totally real or CM field. Let $U\subset V_{r,n}$ be a quadratic form of signature $(2,r-2)$ which admits an $E$-module structure satisfying condition (C) and let $m=\dim_{E}U$.
If $E$ is a totally real field and $m\geq 3$, there is a family of complex projective HK manifolds $X$ with $b_{2}(X)=r$ such that a very general member $X$ has the properties that $T_{X,{\mathbf{Q}}}\simeq U$ as quadratic forms and $E={\rm End}_{\rm Hdg}(T_{X,{\mathbf{Q}}})$. The dimension of this family is $m-2$.
If $E$ is a CM field and $m>1$, there is a family of complex projective HK manifolds $X$ with $b_{2}(X)=r$ such that a very general member $X$ has the properties that $T_{X,{\mathbf{Q}}}\simeq U$ as quadratic forms and $E={\rm End}_{\rm Hdg}(T_{X,{\mathbf{Q}}})$. The dimension of this family is $m-1$.
@@
Proof. In case $E$ is totally real, any $\omega\in\mathbb{P}(U_{\sigma_{0}}\otimes_{\mathbf{R}}\mathbf{C})$ with $q(\omega,\omega)=0$ and $q(\omega,\bar{\omega})>0$ defines a Hodge structure on $U\subset V_{r,n}$ with $U^{2,0}=\mathbf{C}\omega$ which induces one on $V_{r,n}$ and thus on $L_{r,n}$. By the surjectivity of the period map, $L_{r,n}$ is Hodge isometric to $H^{2}(X,{\mathbf{Z}})$ for a HK manifold $X$ (cf. [vGS 23], Proposition 3.7 for the K3 case). Deformations of $\omega\in\mathbb{P}(U_{\sigma_{0}}\otimes_{\mathbf{R}}\mathbf{C})$ induce deformations of $X$. For a very general $\omega$ the Hodge structure on $U$ is simple and there is a Hodge isometry $U\simeq T_{X,{\mathbf{Q}}}$ of K3 type Hodge structures. Notice that $\omega$, a point of an open subset in a quadric in a projective space of dimension $m-1$, depends on $m-2$ parameters.
In case $E$ is a CM field, let $\sigma_{0}$ be an embedding of $E$ in $\mathbf{C}$ for which $(U_{\sigma_{0}}+U_{\bar{\sigma}_{0}})\cap U_{\mathbf{R}}$ has signature $(2,2m-2)$. The eigenspace $U_{\sigma_{0}}$ of $E$ is isotropic for $q$ and a very general $\omega\in\mathbb{P}U_{\sigma_{0}}$ defines a simple polarized Hodge structure on $U$ with $U^{2,0}=\mathbf{C}\omega$. In this case $\omega$ depends on $m-1$ parameters and we obtain an $(m-1)$-dimensional family of HK manifolds whose very general member $X$ has an Hodge isometry $U\simeq T_{X,{\mathbf{Q}}}$.
@@
The HK manifolds deform in a family of dimension $m-1$ such that $E\subset{\rm End}_{\rm Hdg}(T_{\tilde{X}})$ for any member $\tilde{X}$; this should be understood to include the case $m=1$ where the period map has zero-dimensional fibres above the given Hodge structure. Picking a very general member $X_{0}$ of this family (or one of the finitely many in the fiber of the period map in case $m=1$), the claimed equality follows for $X_{0}$ for reason of moduli dimensions. Namely, if $E\subsetneq E^{\prime}={\rm End}_{\rm Hdg}(T_{X_{0},{\mathbf{Q}}})$, then the degree $d^{\prime}$ of $E^{\prime}$ is a multiple of $d$ and with $d^{\prime}m^{\prime}=dm$, we see that $m^{\prime}\leq m/2$, so this can obviously not happen if $m=1$. If $m>1$, then the family, which has CM by $E^{\prime}$, has dimension at most $m/2-1<m-1$. This contradicts the very general choice of $X_{0}$ and thus completes the proof of Theorem 11.2 in the CM case.
If $E$ is totally real of degree $d$ then, since $m\geq 3$, we have an $(m-2)$-dimensional family of HK manifolds $X$ with $E\subset{\rm End}_{\rm Hdg}(T_{X,{\mathbf{Q}}})$. Equality follows again from the dimension of the deformations: if $E\subsetneq E^{\prime}={\rm End}_{\rm Hdg}(T_{X_{0},{\mathbf{Q}}})$ for a very general $X_{0}$, then the degree $d^{\prime}$ of $E^{\prime}$ is a multiple of $d$ and with $d^{\prime}m^{\prime}=dm$ we see that $m^{\prime}\leq m/2$ so the family which has CM or RM by $E^{\prime}$ has dimension at most $m/2-1<m-2$, contradicting the choice of $X_{0}$.
@@
By Lemma 10.1, the members of these families with $\rho>16$ (so that $\rho=18$ or $20$) automatically have CM (cf. [vGS 23, Prop. 7.7] for a similar example).

### v3

@@ 11.1. Hodge structures of K3 type
The polarization is negative definite on $U_{0}:=(U\otimes_{{\mathbf{Z}}}\mathbf{R})\cap U^{1,1}_{X}$ and thus has signature (2,r2)(2,r-2) on $U_{\mathbf{R}}=U_{\mathbf{Z}}\otimes_{\mathbf{Z}}\mathbf{R}$.
A simple K3 type Hodge structure is a member of an r2r-2-dimensional family of K3 type Hodge structures, the general one again being simple, which is parametrized by the choice of an $\omega\in U\otimes_{\mathbf{Z}}\mathbf{C}$ with $q(\omega,\omega)=0$ and $q(\omega,\bar{\omega})>0$; the Hodge structure $U_{\omega}$ determined by $\omega$ has $U^{2,0}_{\omega}=\mathbf{C}\omega$.
Any rational polarized Hodge structure $(U,q)$ of K3 type can be realized geometrically. In fact, it is a Hodge substructure of $H^{2}(KS(U),{\mathbf{Q}})$ where $KS(U)$ is the Kuga Satake abelian variety $KS(U)$ of $U$, which is defined in terms of the Clifford algebra of $(U,q)$ (cf. [vG 00]). Combined with Lefschetz theorems, it follows that these Hodge structures also appear in the second cohomology group of certain projective surfaces. This is quite exceptional, the results on variations of Hodge structures (see [CGG 09]) imply that a general Hodge structure of weight $2$ with $h^{2,0}>1$ or of weight greater than $2$ cannot be a Hodge substructure of a smooth projective variety.
@@ 11.2. Endomorphism algebras of K3 type Hodge structures
$A_{U}\,=\,{\rm End}_{\rm Hdg}(U)\,=\,\{f\in{\rm End}(U):\;f(U^{p,q})\subset U^{p,q}\}\penalty 10000\ .$
Zarhin [Z 83] showed that $A_{U}$ is either a CM field or a totally real field by considering the action of an $f\in A_{U}$ on $U^{2,0}\simeq\mathbf{C}$ and the polarization on $U$. The adjoint of $a\in A_{U}$ for the polarization is the complex conjugate of $a$:
$q(ax,y)\,=\,q(x,\bar{a}y)\penalty 10000\ .$
This adjoint property is equivalent to the existence of an isometry of quadratic forms $U\simeq\mathrm{T}(W)$ for some quadratic (or hermitian) form $W$ over the RM (or CM) field $A_{U}$, see Lemma 5.1. In the RM case, the form $W$ must be such that $W_{\sigma}$ has signature (2,m2)(2,m-2) where $m=\dim W$ and $m\geq 3$ for one embedding and thus $W_{\tau}$ is negative definite for all other embeddings. For totally real fields, there is the following restriction, but for CM fields there is none.
@@
The results in Section 10 characterize the fields $E$, totally real or CM, such that $(U,q)=\mathrm{T}(W)$. The results in Sections 13, 15.1 and 15.2 apply these criteria to specific quadratic forms that arise from Hodge substructures $T\subset H^{2}(X,{\mathbf{Z}})$ of K3 type, where $X$ is a hyperkähler manifold.
@@ 11.3. $K3$ surfaces
We recall the basic facts on K3 surfaces needed for the application of the results obtained in the previous sections. The reader may consult [H 16] for in depth details.
@@
$H^{2}(X,{\mathbf{Z}})\otimes_{{\mathbf{Z}}}\mathbf{C}\,=\,H^{2,0}(X)\oplus H^{1,1}(X)\oplus H^{0,2}(X),\qquad H^{0,2}(X)\,=\,\overline{H^{2,0}(X)}\penalty 10000\ ,$
@@ 11.4. Hyperkähler manifolds
A hyperkähler (HK) manifold, also called irreducible holomorphic symplectic manifold (IHSM), is a simply connected compact complex Kähler manifold $X$ with trivial canonical bundle, see [H 99], [H 12] for comprehensive overviews covering all the prerequisites for this paper.
A 2-dimensional HK manifold is a K3 surface. Currently we know of four higher dimensional families of such manifolds, listed in Table 1 (following [R 08, p. 78])
@@
The second integral cohomology group $H^{2}(X,{\mathbf{Z}})$ of a HK manifold shares many of the properties of the one of a K3 surface. It is a free $\mathbf{Z}$}-module of rank $r=b_{2}(X)$ with an even quadratic form $q$ (not unimodular in general), the Beauville-Bogomolov-Fujiki-form (or BBF-form), of signature (3,r3)(3,r-3). A HK manifold $X$ is projective if and only if $\mathrm{Pic}(X)$ contains an element $D$ with $q(D,D)>0$. The Hodge decomposition for a projective HK manifold has
$H^{2,0}(X)\,=\,\mathbf{C}\omega\;\;\;\text{ and }\;\;\;\mathrm{Pic}(X)=\omega^{\perp}\cap H^{2}(X,{\mathbf{Z}})\penalty 10000\ ,$
where $\omega\in H^{2}(X,\mathbf{C})$ is (the class of) a nowhere zero holomorphic two form. The BBF-form defines a polarization $q$ on the Hodge structure $T_{X}:=\mathrm{Pic}(X)^{\perp}$ and $(T_{X},q)$ is a simple Hodge structure of K3 type. We write $L_{r,n}$ for the lattice $(H^{2}(X,{\mathbf{Z}}),q)$ (which only depends on $r=b_{2}(X)$ and $n=\dim X$):
$L_{r,n}\,\simeq\,(H^{2}(X,{\mathbf{Z}}),q),\quad V_{r,n}\,:=\,L_{r,n}\otimes_{\mathbf{Z}}{\mathbf{Q}}\,\simeq\,(H^{2}(X,{\mathbf{Q}}),q)\penalty 10000\ .$
@@ 11.5. Surjectivity of the period map for HK manifolds
The main result from complex geometry that we need to establish the existence of HK manifolds with RM or CM is the surjectivity of the period map. This is the fact ([H 99, Theorem 8.1]), that if $L_{r,n}=(V_{\mathbf{Z}},q_{V})$ is a lattice which is isometric to $(H^{2}(X,{\mathbf{Z}}),q)$ for a HK manifold $X$ and given an $\omega\in V_{\mathbf{Z}}\otimes_{{\mathbf{Z}}}\mathbf{C}$ with $q_{V}(\omega,\omega)=0$ and $q_{V}(\omega,\overline{\omega})>0$, then there exists a HK manifold $X_{\omega}$ with an isometry $(H^{2}(X_{\omega},{\mathbf{Z}}),q)\simeq(V_{\mathbf{Z}},q)$ that maps $H^{2,0}(X_{\omega})$ to $V^{2,0}:=\mathbf{C}\omega$. In standard terminology, this concerns the image of $\omega$ in the period space
${\mathcal{P}}_{r,n}\,:=\,\{[\omega]\in\mathbb{P}(V_{\mathbf{Z}}\otimes_{\mathbf{Z}}\mathbf{C}):\;q_{V}(\omega,\omega)\,=\,0,\quad q_{V}(\omega,\overline{\omega})>0\}.$
Moreover, if there exists an $h\in V_{\mathbf{Z}}\cap\omega^{\perp}$ with $q_{V}(h,h)>0$, then $X_{\omega}$ is a projective HK manifold. From the Lefschetz $(1,1)$-theorem it follows that
(11.1) $\mathrm{Pic}(X_{\omega})\,=\,\{x\in H^{2}(X_{\omega},{\mathbf{Z}}):\,q(x,\omega)\,=\,0\}.$
@@
For K3 surfaces there is the Torelli theorem which asserts that $X_{\omega}$ is uniquely determined by $\omega$ up to isomorphism. This does not hold in general for HK manifolds, see [H 12].
@@ 12.1. The condition (C)
Let $E$ be a totally real or CM number field of degree $d$. Let $U=(U,q)$ be a quadratic form over $\mathbf{Q}$} of signature (2,r2)(2,r-2).
@@ 12.2. Geometric realizations
Let $(U,q)$ a quadratic form over $\mathbf{Q}$} which is a direct summand of the quadratic form $V_{r,n}\simeq(H^{2}(X,{\mathbf{Q}}),q)$ as in Table 1; notice that this quadratic form only depends on $r=b_{2}(X)$ and $n=\dim X$. Let $E$ be a field such that $U$ admits an $E$-module structure satisfying condition (C). Then the choice of a suitable, general, $\omega$ in a suitable eigenspace of the $E$-action on $U_{\mathbf{C}}$ determines a Hodge structure on $V_{r,n}$ and thus on $L_{r,n}\subset V_{r,n}$. Assuming that we are in one of the four known deformation types from Table 1, this determines a finite number of HK manifolds $X$ by the surjectivity of the period map (and $X$ is unique if $r=22$), together with a Hodge isometry $U\simeq T_{X,{\mathbf{Q}}}$ and $E\subset A_{X}$ (cf. [H 12]); equality can often be established with dimension arguments.
The results obtained in the previous sections on quadratic forms allow us to find examples of such $E$ and $U$. Indeed, for fixed $E$-action on $V_{r,n}=L_{r,n}\otimes_{\mathbf{Z}}{\mathbf{Q}}$, the periods $\omega]\in{\mathcal{P}}_{r,n}$ that define HK manifolds $X$ with RM resp. CM by $E$ are those in a complex submanifold of the period space. (to be made explicit in the proof of Thm 12.2). We refer to the dimension of the submanifold as the dimension of the family of HK manifolds of dimension $2n$ and $b_{2}(X)=r$.
With this in mind, the following result provides the key geometric ingredient to prove Theorems A and B.
@@
Let $V_{r,n}=L_{r,n}\otimes_{\mathbf{Z}}{\mathbf{Q}}$ where $L_{r,n}$ is a lattice from Table 1. Let $E$ be a totally real or CM field. Let $U\subset V_{r,n}$ be a quadratic form of signature (2,r2)(2,r-2) which admits an $E$-module structure satisfying condition (C) and let $m=\dim_{E}U$.
If $E$ is a totally real field and $m\geq 3$, there is a family of complex projective HK manifolds $X$ with $b_{2}(X)=r$ such that a very general member $X$ has the properties that $T_{X,{\mathbf{Q}}}\simeq U$ as quadratic forms and $E={\rm End}_{\rm Hdg}(T_{X,{\mathbf{Q}}})$. The dimension of this family is m2m-2.
If $E$ is a CM field and $m>1$, there is a family of complex projective HK manifolds $X$ with $b_{2}(X)=r$ such that a very general member $X$ has the properties that $T_{X,{\mathbf{Q}}}\simeq U$ as quadratic forms and $E={\rm End}_{\rm Hdg}(T_{X,{\mathbf{Q}}})$. The dimension of this family is m1m-1.
@@
Proof. In case $E$ is totally real, any $\omega\in\mathbb{P}(U_{\sigma_{0}}\otimes_{\mathbf{R}}\mathbf{C})$ with $q(\omega,\omega)=0$ and $q(\omega,\bar{\omega})>0$ defines a Hodge structure on $U\subset V_{r,n}$ with $U^{2,0}=\mathbf{C}\omega$ which induces one on $V_{r,n}$ and thus on $L_{r,n}$. By the surjectivity of the period map, $L_{r,n}$ is Hodge isometric to $H^{2}(X,{\mathbf{Z}})$ for a HK manifold $X$ (cf. [vGS 25], Proposition 3.7 for the K3 case). Deformations of $\omega\in\mathbb{P}(U_{\sigma_{0}}\otimes_{\mathbf{R}}\mathbf{C})$ induce deformations of $X$. For a very general $\omega$ the Hodge structure on $U$ is simple and there is a Hodge isometry $U\simeq T_{X,{\mathbf{Q}}}$ of K3 type Hodge structures. Notice that $\omega$, a point of an open subset in a quadric in a projective space of dimension m1m-1, depends on m2m-2 parameters.
In case $E$ is a CM field, let $\sigma_{0}$ be an embedding of $E$ in $\mathbf{C}$ for which $(U_{\sigma_{0}}+U_{\bar{\sigma}_{0}})\cap U_{\mathbf{R}}$ has signature (2,2m2)(2,2m-2). The eigenspace $U_{\sigma_{0}}$ of $E$ is isotropic for $q$ and a very general $\omega\in\mathbb{P}U_{\sigma_{0}}$ defines a simple polarized Hodge structure on $U$ with $U^{2,0}=\mathbf{C}\omega$. In this case $\omega$ depends on m1m-1 parameters and we obtain an (m1)(m-1)-dimensional family of HK manifolds whose very general member $X$ has an Hodge isometry $U\simeq T_{X,{\mathbf{Q}}}$.
@@
[…]
If $E$ is totally real of degree $d$ then, since $m\geq 3$, we have an (m2)(m-2)-dimensional family of HK manifolds $X$ with $E\subset{\rm End}_{\rm Hdg}(T_{X,{\mathbf{Q}}})$. Equality follows again from the dimension of the deformations: if $E\subsetneq E^{\prime}={\rm End}_{\rm Hdg}(T_{X_{0},{\mathbf{Q}}})$ for a very general $X_{0}$, then the degree $d^{\prime}$ of $E^{\prime}$ is a multiple of $d$ and with $d^{\prime}m^{\prime}=dm$ we see that $m^{\prime}\leq m/2$ so the family which has CM or RM by $E^{\prime}$ has dimension at most m/21<m2m/2-1<m-2, contradicting the choice of $X_{0}$.
@@
By Lemma 11.1, the members of these families with $\rho>16$ (so that $\rho=18$ or 2020) automatically have CM (cf. [vGS 25, Prop. 7.7] for a similar example).
Remark 12.6.
There are also applications to Fano varieties with K3-type Hodge structures. For instance, the Fano variety of lines $F(X)$ on a cubic fourfold $X$ is a HK variety, and by [BD 85] there is a correspondence on the product $X\times F(X)$ which induces an isomorphism on the Hodge structures. This translates RM or CM structures from one to the other.